OpenStudy (anonymous):

A taxi charges a fee of $5 per ride plus $0.30 per mile. Write an algebraic expression for the cost of m miles. What is the cost for 5 miles? (1 point) 0.30m + 5; $6.50 0.30m – 5; $1.50 0.30m + 5; $1.50 0.30m – 5; $6.50

pooja195 (pooja195):

Fixed fee: 5 1 mile: $0.30 x 1 2 miles: $0.30 x 2 3 miles: $0.30 x 3 m miles: $0.30 x m = 0.30m Total cost= fixed fee + per-mile fee Cost = 0.30m + 5 For 5 miles: Cost = 0.30(5) + 5 Cost = 1.50 + 5 Cost = 6.50 first one (someone elses work)

OpenStudy (anonymous):

a. , b. , c. , d.

pooja195 (pooja195):

a